OpenStudy (nostalgia):

Vivian is making a quilt and she has determined she needs 870 square inches of orange fabric and 1417 square inches of green. How many square yards of each material will she need? Round your answers up to the nearest quarter yard. Ive done the math multiple times and i seem to be doing somthing wrong it says to round to the nearest quarter yard which ive done and it still says its incorrect, anyone help out?

OpenStudy (misty1212):

HI!!

OpenStudy (misty1212):

guess you have to convert square inches to square yards, right?

OpenStudy (misty1212):

kind of a pain actually do you know how to do it?

OpenStudy (nostalgia):

I do not, i figured that you have to divide twice but i dont think thats right.

OpenStudy (misty1212):

oh no

OpenStudy (misty1212):

there are 12 inches in a foot, 3 feet in a yard, so 36 inches in a yard, so so one square yard is \(36^2\) inches

OpenStudy (misty1212):

in other words, you have to divide each of those numbers by \(36^2=1296\)

OpenStudy (nostalgia):

That makes alot more sense then what i was trying...

OpenStudy (misty1212):

so the first is \[870\div 1296=0.671296\]

OpenStudy (misty1212):

to the nearest quarter? i guess \(\frac{3}{4}\) is the closest, i.e. \(0.75\)

OpenStudy (nostalgia):

Thats where i was wrong! See i got that answer as well but i wasnt quite sure what the whole rounding to the nearest quarter, I got that correct finally so can you explain a little what rounding to the nearest quarter is then real quick? thats my problem

OpenStudy (misty1212):

rounding to the nearest quarter, in this case i guess you need decimals \[\frac{1}{4}=0.25, \frac{1}{2}=0.5,\frac{3}{4}=.75\] etc

OpenStudy (misty1212):

\[1417\div 1296\] probably rounds to \(1\)

OpenStudy (misty1212):

ooh no it says round UP

OpenStudy (misty1212):

so i guess \(1.25\) or one and a quarter

OpenStudy (nostalgia):

Yea i got it, thank you so much stuck on this question for so long you have no idea thank you! (: